9. The following circuit operates only if there is a path of functional devices from left to right. If 0.9, what is the components work independently of one another and P(component works) probability that the circuit operates?
Answer #1

P(A-B) works =1-P(none of A and B works) =1-(1-0.9)*(1-0.9)=0.99

P(CD and EF works)=1-P(none of CD and EF works)=1-(1-0.9*0.9)*(1-0.9*0.9)=0.9639

hence P(circuit operates)=P(AB works)*P(CD and EF works)*P(G works)=0.99*0.9639*0.9=0.858835